# Efficient Prompting for LLM-based Generative Internet of Things

### v2

@@ V-B2 The impact of quantization
| LLM | Parameter Precision | EM Acc |
| Mixtral-8x7B | 16-bit | 63.33 |
| 8-bit | 62.20 | |
| 4-bit | 60.52 |
@@ VI Conclusion and Future Work
In this study, we propose a LLM-based GIoT system, which can be deployed in a local network setting to address the security concerns of many scenarios. To demonstrate the proposed LLM-based GIoT system, we use a challenging semi-structured Table-QA problem as a case study and propose a three-stage prompting solution to alleviate the issues caused by complex structures, heterogeneous data types, huge tables, and LLMs’ limitations. The proposed prompting solution uses a statistics table in the first stage and sub-tables in the following stages, which can reduce the inference cost and improve the performance when the original table is huge. We define a series of atomic operations to guide the demonstration crafting and selection, which can reduce reasoning errors. Besides, we use the task-correction stage to correct the failure code caused by the heterogeneous data types and use a default answer as the final answer when the generated Python code fails to run even after the task-correction step, which can be caused by the complex structure or the limitations of the LLM. As demonstrated in Section V, designing tailored prompting methods can improve the performance of open-source LLMs, achieving state-of-the-art performance, and the proposed LLM-based GIoT system can be easily extended by adding task-specific prompt instructions and demonstrations to the system. Besides, as the proposed GIoT system is designed to deploy in a edge server, applying quantization to the LLM can be a good option to reduce the hardware requirements, as discussed in Section V-B2. In this study, we focus on text data, while IoT devices can generate data in multiple data types, such as time series and images. Therefore, extending the current system to handle data in various modalities can be a further direction.
@@ Acknowledgement
This work is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) under the CREATE TRAVERSAL and DISCOVERY programs.

### v3

@@ V-B2 The impact of quantization
#GPU | Model | Parameter | EM | AVG Prompt | AVG Completion |
Size | Precision | Acc | Throughput | Throughput | |
4 | 87G | 16-bit | 63.33 | 743.8 tokens/s | 58.8 tokens/s |
4 | 48G | 8-bit | 62.20 | 553.7 tokens/s | 39.7 tokens/s |
2 | 48G | 8-bit | 62.20 | 377.6 tokens/s | 26.5 tokens/s |
4 | 25G | 4-bit | 60.52 | 581.3 tokens/s | 49.3 tokens/s |
2 | 25G | 4-bit | 60.52 | 427.7 tokens/s | 35.9 tokens/s |
1 | 25G | 4-bit | 60.52 | 286.8 tokens/s | 24.3 tokens/s |
@@ V-B4 Analysis on inference cost
As mentioned in Section V-A, we deployed the LLMs with vLLM [59] on a workstation with 8 NVIDIA A100 40G GPUs. We utilized a single GPU for the Mistral-7B and DeepSeek-7B and four GPUs for the Mixtral-8x7B and DeepSeek-67B. The throughput performances of each model (with respect to prompt and completion) are summarized in Table VIII. As shown in Table VIII, the throughput performance of the DeepSeek-67B is significantly lower than other models for both prompt and completion, which is caused by the overhead of using tensor parallelism. Specifically, we observed that the number of components of the DeepSeek-67B deployed in the CPUs is much larger than that of Mixtral-8x7B for the tensor parallelism, making it much slower than Mixtral-8x7B which also needs to be deployed using tensor parallelism. By contrast, the Mistral-7B and DeepSeek-7B, both of which can be deployed in a single GPU, can show similar throughput performances. At last, it is worth mentioning that the inference overheads caused by the Prompt Management Module and the Post-processing Module are negligible, because they can be finished in milliseconds for our cases, while the LLM inference usually needs at least a few seconds.
Model | #GPU | AVG Prompt | AVG Completion |
Throughput | Throughput | ||
Mixtral-8x7B | 4 | 743.8 tokens/s | 58.8 tokens/s |
Mistral-7B | 1 | 633.0 tokens/s | 57.7 tokens/s |
DeepSeek-67B | 4 | 320.7 tokens/s | 20.1 tokens/s |
DeepSeek-7B | 1 | 705.2 tokens/s | 59.4 tokens/s |
@@ VI Conclusion and Future Work
In this study, we propose an LLM-based GIoT system, which can be deployed in a local network setting to address the security concerns of many scenarios. To demonstrate the proposed LLM-based GIoT system, we use a challenging semi-structured Table-QA problem as a case study and propose a three-stage prompting solution to alleviate the issues caused by complex structures, heterogeneous data types, huge tables, and LLMs’ limitations. The proposed prompting solution uses a statistics table in the first stage and sub-tables in the following stages, which can reduce the inference cost and improve the performance when the original table is huge. We define a series of atomic operations to guide the demonstration crafting and selection, which can reduce reasoning errors. Besides, we use the task-correction stage to correct the failure code caused by the heterogeneous data types and use a default answer as the final answer when the generated Python code fails to run even after the task-correction step, which can be caused by the complex structure or the limitations of the LLM. As demonstrated in Section V, designing tailored prompting methods can improve the performance of open-source LLMs, achieving state-of-the-art performance, and the proposed LLM-based GIoT system can be easily extended by adding task-specific prompt instructions and demonstrations to the system. Besides, as the proposed GIoT system is designed to deploy in an edge server, applying quantization to the LLM can be a good option to reduce the hardware requirements, as discussed in Section V-B2. In this study, we focus on text data, while IoT devices can generate data in multiple data types, such as time series and images. Therefore, extending the current system to handle data in various modalities can be a further direction. Besides extending the system to multiple modalities, improving computational efficiency and reducing hardware requirements are also critical to the system. Therefore, the caching mechanisms which caches intermediate results or frequently used outputs, model pruning which can reduce the size of LLM models, and many others are also promising directions.
@@ Acknowledgement
This work is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) under the CREATE TRAVERSAL Program and NSERC DISCOVERY program, and also in part by the National Research Foundation, Singapore, and Infocomm Media Development Authority under its Future Communications Research & Development Programme, AI Singapore Programme (FCP-NTU-RG-2022-010 and FCP-ASTAR-TG-2022-003), Singapore Ministry of Education (MOE) Tier 1 (RG87/22), and the NTU Centre for Computational Technologies in Finance (NTU-CCTF).